# Enabling Fast and Accurate Neutral Atom Readout through Image Denoising

### v1

@@ Abstract.
Neutral atom quantum computers hold promise for scaling up to hundreds of thousands of qubits, but their progress is constrained by slow qubit readout. Measuring qubits currently takes milliseconds—much longer than the underlying quantum gate operations—making readout the primary bottleneck in deploying quantum error correction. Because each round of QEC depends on measurement, long readout times increase cycle duration and slow down program execution.
@@
We show that image denoising can resolve this tension. Our framework, GANDALF, uses explicit denoising using image translation to reconstruct clear signals from short, low-photon measurements, enabling reliable classification at up to shorter readout times. Combined with lightweight classifiers and a pipelined readout design, our approach both reduces logical error rate by up to and overall QEC cycle time up to compared to state-of-the-art CNN-based readout for Cesium (Cs) Neutral Atom arrays.
@@ 1. Introduction
Neutral atom quantum computers have emerged as a promising platform for building scalable quantum systems. In recent years, they have demonstrated the ability to trap and control thousands of individual atoms—with systems reaching up to 3,000 qubits (Chiu et al., 2025). These systems are also making progress towards fault tolerance (Sales Rodriguez et al., 2025), with successful demonstrations of quantum error correction (QEC)(Bluvstein et al., 2024), and are experimentally demonstrating applications in areas such as analog quantum simulation(Bernien et al., 2017), and optimization (Graham et al., 2022).
Despite these advancements, a major challenge for neutral atom platforms is their relatively slow operation speeds. Table 1 summarizes the typical durations of key operations—including two-qubit gates, single-qubit gates, readout, and full QEC cycles for leading quantum error correction codes—for systems from QuEra and various academic labs. For reference, we also show the operational latency for Google’s superconducting hardware.
@@
Why is readout slow in neutral atom architectures? In neutral atom systems, all atoms in the array are illuminated with a laser. If a qubit is in state , it will scatter photons; if it is in state , it will remain dark. These scattered photons are collected by a highly sensitive camera to determine each qubit’s state. However, each atom emits only a small number of photons, and the optics can only collect a small fraction of them.
@@
| 34 ns | 500 ns | 0.5 | |
| QuEra | 400 ns | 10 ms | 0.75 |
@@
To measure a qubit accurately, the system needs to collect more photons from the atom. It can do this by either increasing the laser power or extending the exposure time. Increasing the power risks overheating the atom, which can cause it to escape its trap, or jump to an unwanted energy level, or experience crosstalk. The other option is extending the exposure to collect more photons, but this slows down readout. As a result, the system faces a trade-off: faster measurements collect fewer photons and are less reliable, while slower measurements improve accuracy but take more time.
@@
Ideally, we want fast readout without using high laser power, which translates to doing classification at significantly low photon counts. Most existing readout classifiers (Phuttitarn et al., 2024) rely on high photon counts to distinguish between qubit states. Higher photon counts create a clear separation between and states. But this assumption breaks down for fast, low photon count measurements, where classical classifiers struggle to maintain the same level of accuracy. This limitation presents a significant bottleneck for enhancing QEC, calibration, and application-level performance.
@@
In this work, we demonstrate that denoising can enable high-fidelity readout even at low photon counts and can be effectively used for accelerating the readout for neutral atom quantum computers. We utilize a dataset of paired fluorescence images from a Cesium (Cs) atom array, where a high Signal to Noise Ratio (SNR) path provides long-exposure ground truth labels and an attenuated path produces noisy, short-exposure measurements (Phuttitarn et al., 2024). We observe that conventional CNN classifiers, effective in natural-image domains experience degraded performance because bright and dark histograms overlap and neutral atom images lack semantic richness. This motivates a different approach of processing the noisy images prior to classification to improve the SNR by filtering the noise, instead of classifying them directly.
@@
Our central design insight is to decouple signal recovery from classification. We introduce GANDALF, a generative adversarial framework that denoises low-photon images, restoring per-site fidelity and suppressing photon noise and crosstalk. By training on paired datasets of low- and high-photon fluorescence frames from cesium atom arrays (Phuttitarn et al., 2024), our approach reconstructs high-fidelity images from fast, low-photon measurements—enabling rapid and accurate qubit state classification. Figure 1 shows that GAN can significantly improve the signal-to-noise ratio for a short-exposure atom array image, enabling high-fidelity readout even at low photon counts.
@@
Moreover, with denoising, the classification task becomes far easier, enabling the use of lightweight site-specific models such as feedforward networks or matched filter hybrids. These compact classifiers match the accuracy of larger CNNs while cutting latency by up to five times and parameter count by more than an order of magnitude, making them practical for arrays with thousands of qubits. Equally important, we propose the readout pipeline to overlap classification with image acquisition. By pipelining these stages, classification adds only once to total latency rather than every cycle, eliminating it as a recurring bottleneck as shown in Figure 1(b). Pipelining of readout denoising and classification enables room for using sophisticated ML models that can maintain high accuracy with low SNR to reduce readout time.
Finally, by making the denoising model fully convolutional, the same network trained on small calibration arrays seamlessly generalizes to larger lattices, maintaining per-site inference cost nearly constant as system size grows. This is crucial as generating fully occupied qubit lattices can be slow experimentally due to slow loading.
@@
Our evaluations show that this combination of denoising GAN and a lightweight classifier, such as a shallow Feed Forward Neural Network (FNN), delivers significant gains. At exposures, GANDALF reduces readout error by compared to CNN baselines. The dataset we use is imperfect, and we expect higher-quality data to be available in the near future. Therefore, to understand the best-case improvements with denoising, we create a dataset generated via confidence-based filtering, and on this smaller post-selected dataset, we observe sub-1% inaccuracy even at the shortest duration of . This highlights the effectiveness of our approach in reducing readout duration to the sub-millisecond region for Cs atom arrays. At the system level, logical error rates improve dramatically—up to reduction for bivariate bicycle codes and for surface codes—demonstrating that faster, denoised readout directly strengthens fault tolerance. Lightweight classifiers achieve inference in the 100s range, and pipelining shortens overall QEC cycle time by , reducing idling-induced decoherence. Even with commodity GPUs with 4K CUDA cores, the denoiser scales from to and even larger arrays with sublinear growth in latency, ensuring feasibility for next-generation processors. This paper makes the following contributions:
@@
We demonstrate denoising enables reliable state discrimination at much shorter exposure times, making low-photon and pipelined readouts viable for fault-tolerant execution.
By adapting generative networks to the unique structure of fluorescence snapshots, we show GAN-based denoising outperforms conventional CNNs in the low-SNR regime.
Our fully convolutional framework maintains per-site fidelity while effectively suppressing noise and crosstalk, and it generalizes robustly from small calibration arrays to large-scale atom lattices.
@@
GANDALF reduces readout error by up to , lowers logical error rates up to for surface code and up to for BB code, and shortens QEC cycle time by , directly addressing readout bottleneck in neutral atom quantum computers.
@@ 2. Background
2.1. Neutral atom quantum computing
Neutral atom platforms have emerged as a promising architecture for scalable quantum computing, offering flexible qubit layouts, long coherence times, and high connectivity, (Wintersperger et al., 2023b; Saffman, 2019). In these systems, individual atoms—typically species such as Rubidium (Rb), Cesium (Cs), Strontium (Sr), or Ytterbium (Yb)—serve as qubits, with internal electronic or hyperfine states encoding the logical and states.
Qubits are held in space using optical tweezers: tightly focused laser beams that create potential wells to trap single atoms. Arrays of tweezers are generated using acousto-optic deflectors or spatial light modulators, allowing dynamic reconfiguration and parallel operation. The vacuum environment and tight localization of atoms enable long coherence times, while high-numerical-aperture lenses allow for high-fidelity optical control and measurement.
Single-qubit gates are implemented using resonant laser or microwave pulses that drive coherent rotations between the two qubit states. Recent experiments (Madjarov et al., 2020; Wintersperger et al., 2023b) performed these operations in parallel across the array, with gate fidelities exceeding 99.5%. Two-qubit entangling gates are typically realized via the Rydberg blockade mechanism: atoms are excited to high-energy Rydberg states where strong dipole-dipole interactions prevent simultaneous excitation of nearby qubits (Isenhower et al., 2010; Levine et al., 2019). This interaction enables fast, high-fidelity two-qubit gates with interaction ranges spanning several micrometers, making neutral atom systems especially well-suited for two-dimensional lattice-based quantum error correction codes and programmable quantum simulations.
@@ 2.2. Readout in neutral atom architecture
In neutral atom quantum computers, readout relies on detecting whether an atom is in a particular internal state. These qubit states are typically encoded in long-lived hyperfine or electronic levels, and readout works by shining a laser tuned to excite only one of the qubit states. If the atom is in the excited state (e.g., ), it will scatter photons and appear bright under a camera. If it is in the other state (e.g., ), it remains dark. By capturing this fluorescence with a sensitive camera, the system determines the atom’s state.
Although conceptually simple, this readout process is slow compared to the speed of quantum gates. To confidently distinguish between bright and dark states, the system must collect enough photons to overcome background noise and random fluctuations. This requires either using a stronger laser or increasing the exposure time. A stronger laser can cause heating, which may knock the atom out of its trap or excite it to unwanted energy levels. On the other hand, a longer exposure slows down the readout and limits how quickly quantum circuits can be executed.
Additionally, the fidelity of readout depends on many experimental parameters, including the numerical aperture of the imaging system, the qubit species, and spacing between atoms. For example, atoms placed closely together can interfere with each other, a problem known as crosstalk. To mitigate these issues, researchers are exploring advanced methods such as machine learning and denoising techniques to improve readout fidelity at shorter measurement times.
Overall, while readout is essential for extracting useful information from a quantum computer, its speed and reliability pose a major challenge for scaling up neutral atom systems and supporting fast feedback in quantum error correction.
@@ 2.3. Why does fast readout matter?
The execution time of a fault-tolerant quantum program can be coarsely approximated as:
Each component in this model encapsulates a key bottleneck to scalable, fault-tolerant quantum computation:
Non-Clifford Gates (e.g., gates) are necessary for universal quantum computation. They are the slowest operations due to magic state distillation (Gidney, 2025).
-Rounds represent the number of QEC iterations needed to suppress logical errors for a code of distance . As expected, larger values are required for lower physical error rates, increasing overall runtime.
QEC cycle time is the duration of one round of quantum error correction, where errors are detected and corrected to keep logical qubits free of errors.
@@
Atom loading and rearrangement overhead.
In neutral atom platforms, building a defect-free 2D qubit lattice requires stochastic loading followed by iterative rearrangement. This initialization process is slow, imaging intensive, and requiring repeated fluorescence-based measurements to identify occupied sites and reposition atoms using optical tweezers. Fast, high-fidelity readout accelerates each stage of this pipeline—enabling denser lattices, reducing the probability of atom loss during relocation, and improving experiment yield. Though often hidden behind “setup time,” inefficiencies here can dominate runtime in large-scale systems or in hybrid pipelines where reloading is triggered adaptively.
QEC bootstrapping and syndrome verification.
Initialization is not limited to loading alone. For QEC codes to start correctly, ancilla qubits must be verified through measurement. Faster readout enables quicker feedback, reducing dead time and improving the throughput of syndrome generation across the code lattice.
Pipelining and reuse of physical infrastructure.
In principle, multiple Magneto-Optical Traps (MOTs) can be deployed in parallel to mask atom loading time, but doing so increases system complexity, cost, and physical footprint. Faster readout alleviates the need for deep pipelining by making each MOT pass more productive, reducing the amortized cost per qubit and easing scalability constraints.
While readout speed directly governs QEC cycle time, its influence is far broader—impacting lattice preparation, code startup, and system utilization efficiency. Fast readout is not just a lever for shorter execution times; it is a key enabler for practical, reliable, and reconfigurable fault-tolerant quantum computing on neutral atom platforms.
3. Challenges in Neutral Atom Readout
3.1. State of the art
Qubit readout remains a key bottleneck in neutral atom quantum computing, with significant variations in fidelity and speed across available atomic species, including Cesium (Cs), Rubidium (Rb), Ytterbium (Yb), and Strontium (Sr) illustrated in Figure 2. The performance of fluorescence-based state detection depends strongly on species-specific properties such as cycling transition strength, background scattering, and susceptibility to state leakage. For instance, alkaline-earth(-like) atoms such as Yb and Sr offer narrow-line transitions that support high-fidelity imaging with minimal atom loss (Saskin et al., 2019; Madjarov et al., 2020), but they often require cryogenic cooling and complex laser systems. In contrast, alkali atoms like Cs and Rb are experimentally simpler to trap and manipulate, but their readout typically involves trade-offs between speed, fidelity, and trap loss due to heating and off-resonant scattering (Endres et al., 2013; Phuttitarn et al., 2024). Improving readout performance across all species is crucial for supporting quantum error correction, which relies on fast and repeated measurements of ancilla qubits. Each atomic species imposes distinct system-level constraints such as photon budget, trap spacing, and detection optics necessitating a careful balance between fidelity, speed, hardware complexity, and scalability.

### v2

@@ Abstract.
Neutral atom quantum computers hold promise for scaling up to hundreds of thousands or more qubits, but their progress is constrained by slow qubit readout. Parallel measurement of qubit arrays currently takes milliseconds,—much longer than the underlying quantum gate operations—making readout the primary bottleneck in deploying quantum error correction. Because each round of QEC depends on measurement, long readout times increase cycle duration and slow down program execution.
@@
We show that image denoising can resolve this tension. Our framework, GANDALF, uses explicit denoising using image translation to reconstruct clear signals from short, low-photon measurements, enabling reliable classification at up to shorter readout times. Combined with lightweight classifiers and a pipelined readout design, our approach both reduces logical error rate by up to and overall QEC cycle time up to compared to state-of-the-art convolutional neural network (CNN)-based readout for Cesium (Cs) Neutral Atom arrays.
@@ 1. Introduction
Neutral atoms quantum computers have emerged as a promising platform for building scalable quantum computers. In recent years, they have demonstrated the ability to trap and control thousands of individual atoms—with systems reaching up to 6,100 qubits (Manetsch_2025; Chiu_2025). These systems are also making progress towards fault tolerance (sales2025experimental; Bedalov2024), with successful demonstrations of Quantum Error Correction (QEC)(bluvstein2024logical), and are experimentally demonstrating applications in areas such as analog quantum simulation(bernien2017probing), and optimization (graham2022multi).
Despite these advancements, a major challenge for neutral atom platforms is their relatively slow operation speeds. Table 1 summarizes the bottlenecks in QEC cycle—including two-qubit gates, readout, and fraction of readout in full quantum error correction cycles for leading quantum error correction codes—for the neutral atom systems (Wintersperger_2023). For reference, we also show the operational latency for Google’s superconducting hardware (google2025).
@@
Why is readout slow in neutral atom architectures? In neutral atom systems, all atoms in the array are illuminated with a laser. If a qubit is in state , it will emit photons; if it is in state , it will remain dark. These scattered photons are collected by a highly sensitive camera to determine each qubit’s state. However, each atom can at most emit a fixed number of photons per unit time, and the optics can only collect and detect a small fraction of the emitted photons.
@@
| Google (google2025) | 34 ns | 500 ns | 0.5 |
| Neutral Atom (Wintersperger_2023) | 400 ns | 1 ms | 0.75 |
@@
To measure a qubit accurately, the system needs to collect a sizable number of photons from the atom. To increase the collected photons, we can increase the laser power or increase the duration of the readout pulse. Increasing the power only increases the photon emission rate up to a limit set by atomic parameters, and risks overheating the atom, which can cause it to escape its trap, or jump to an unwanted energy level, or experience crosstalk. The other option is to stretch the readout pulse, increasing exposure time to collect more photons, which slows down readout. As a result, designers face a trade-off: faster readout collect fewer photons and are less reliable, while slower readout improve accuracy but take more time.
@@
Ideally, we want fast readout without using high laser power, which translates to performing classification at significantly lower photon counts. Most existing readout classifiers (phuttitarn2023enhanced; kent2025efficientmeasurementneutralatomqubits) rely on high photon counts to distinguish between qubit states. Higher photon counts create a clear separation between and states. But this assumption breaks down for fast, low-photon-count measurements, because it’s hard to distinguish state-dependent illumination of an atomic qubit from background noise caused by spurious photons.
We observe that the existing classifiers struggle to maintain the same level of accuracy at low photon count due to low signal to noise (SNR) ratio. This limitation presents a significant bottleneck for enhancing quantum error correction, calibration, and application-level performance.
@@
In this work, we demonstrate that denoising can enable high-fidelity readout at low photon counts and can be effectively used for accelerating the readout for neutral atom quantum computers. We utilize a dataset of paired fluorescence images from a Cesium (Cs) atom array, where a high Signal to Noise Ratio (SNR) path provides long-exposure ground truth labels and an attenuated path produces noisy, short-exposure measurements (phuttitarn2023enhanced). We observe that conventional CNN classifiers, at low photon counts, can not differentiate the background noise from qubit-emitting photons. We believe neutral-atom images lack semantic richness of natural images, where typical CNN based computer vision model have shown significant accuracy even under noisy conditions. This motivates a different approach of processing the noisy images prior to classification to improve the SNR by filtering the noise, instead of classifying them directly.
@@
Our central design insight is to decouple signal recovery from classification. We introduce a Generative Adversarial Network for Detection at Low Fluorescence levels (GANDALF), a generative adversarial framework that denoises low-photon images, restoring per-site fidelity and suppressing photon noise and crosstalk improving SNR. By training on paired datasets of low- and high-photon fluorescence frames from cesium atom arrays (phuttitarn2023enhanced), our approach reconstructs high-fidelity images from fast, low-photon measurements—enabling rapid and accurate qubit state classification. Figure 1 shows that GAN can significantly improve the signal-to-noise ratio for a short-exposure atom array image, enabling high-fidelity readout even at low photon counts.
@@
Moreover, with denoising, amplifying the signal, the classification task becomes far easier, enabling the use of lightweight site-specific classifier models such as feedforward networks, matched-filter networks, or their hybrids (kent2025efficientmeasurementneutralatomqubits). These compact classifiers match the accuracy of larger CNNs (phuttitarn2023enhanced) while cutting latency by up to five times and parameter count by more than an order of magnitude, making them practical for arrays with thousands of qubits. Furthermore, we propose a readout architecture with a pipelined denoising and classification, so that they are not in the critical path. This opens the door to using sophisticated ML denoising and classification algorithms running on classical accelerators (nvidia_nvqlink).
@@
Finally, by making the denoising model fully convolutional, the same network trained on small calibration arrays seamlessly generalizes to larger lattices, maintaining a nearly constant per-site inference cost as the system size grows. This is crucial as generating training data for fully occupied qubit lattices can be slow experimentally due to slow loading of atoms on current systems. Our evaluations show that this combination of denoising with GAN and a lightweight classifier, such as a shallow Feed Forward Neural Network (FNN), delivers significant gains. At exposure times of , GANDALF reduces readout error by compared to CNN baselines (phuttitarn2023enhanced). The dataset we use is imperfect, and we expect higher-quality data to be available in the near future. Therefore, to understand the best-case improvements bu using denoising, we create a dataset generated via confidence-based filtering and observe sub-1% inaccuracy even at the shortest duration of . This highlights the effectiveness of our approach in reducing readout duration close to millisecond region for Cesium (Cs) atom arrays.
At the system level, logical error rates improve dramatically—up to reduction for bivariate bicycle codes and for surface codes—demonstrating that faster, denoised readout directly strengthens fault tolerance. Lightweight classifiers achieve inference in the 100 s range, and pipelining shortens overall QEC cycle time by , reducing idling-induced decoherence. Even with commodity GPUs with only 4096 CUDA cores, the denoiser scales from to and beyond, with sublinear latency growth, ensuring scalability for next-generation atom-array processors.
Our work is the first to overcome the low-SNR barrier in neutral-atom readout by reconstructing high-SNR readout images prior to classification. This architectural change shifts the burden of readout accuracy from hardware factors, such as exposure time and laser power, to an upstream denoising stage, enabling accurate, fast readout and substantially lower logical error rates, as well as shorter QEC cycle times. Our design scales naturally to large atom arrays and integrates seamlessly into efforts on connecting neutral atom quantum platforms with GPUs for scaling QEC (nvidia_nvqlink; caldwell2025platform). This paper makes the following contributions:
@@
We demonstrate that denoising enables reliable state discrimination at much shorter exposure times, making low-photon and fast readout viable for fault-tolerant execution.
We show that generative denoising adapted to the non-semantic, site-local structure of neutral-atom readout images yields higher classification accuracy than existing downstream methods operating directly on raw readout images.
We demonstrate scalability with our fully convolutional denoiser whose learned filters apply consistently from small arrays to large atom lattices without architectural changes, maintaining per-atom-site readout accuracy.
@@
We show GANDALF reduces readout error by up to , lowers logical error rates up to for surface code and for the bivariate-bicycle (BB) code, and shortens quantum error correction cycle time by , directly addressing readout bottlenecks in neutral atom quantum computers.
@@ 2. Background
2.1. Qubits and gates for neutral atom quantum computing
Neutral atom platforms have emerged as a promising architecture for scalable quantum computing, offering flexible qubit layouts, long coherence times, and high connectivity (Wintersperger_2023; saffman2019quantum). In these systems, individual atoms—typically species such as Rubidium (Rb), Cesium (Cs), Strontium (Sr), or Ytterbium (Yb)—serve as qubits, with internal states encoding the computational and states.
Qubits are held in space using optical tweezers111Arrays of optical tweezers are generated using acousto-optic deflectors (AOD) or spatial light modulators (SLM), allowing dynamic reconfiguration and parallel operation., the tightly focused laser beams, that create potential wells to trap single atoms. The vacuum environment and tight localization of atoms enable long coherence times, while the high-numerical-aperture lenses allow for the precise optical control and measurement. Single-qubit gates are implemented222In recent experiments (madjarov2020highfidelity; Wintersperger_2023), these operations are performed in parallel across the array, with gate fidelities exceeding 99.5%. using resonant laser or microwave pulses that drive coherent rotations between the two qubit states. Two-qubit entangling gates are typically realized via the Rydberg blockade mechanism333Atoms are excited to high-energy Rydberg states where strong dipole-dipole interactions prevent simultaneous excitation of nearby qubits. which prevents simultaneous excitation of nearby qubits (isenhower2010demonstration; levine2019parallel).
These atomic interactions enable fast, high-fidelity two-qubit gates with interaction ranges spanning several microns, making neutral atom systems especially well-suited for two-dimensional lattice-based quantum error correction codes and programmable quantum simulations.
@@ 2.2. Readout in neutral atom architecture
In neutral atom quantum computers, qubit states are typically encoded in long-lived hyperfine or electronic levels and readout relies on detecting whether the atom is in a particular internal state. Readout works by repeatedly shining a laser tuned to excite only one of the qubit states. If the atom is in the excited state (e.g., ), it will scatter photons and appear bright under a camera. If it is in the other computational state (e.g., ), it remains dark. By capturing this fluorescence with a sensitive camera, the system determines the atom’s state.
Although conceptually simple, this readout process is slow compared to the speed of quantum gates. To confidently distinguish between bright and dark states, the system must collect enough photons to overcome background and camera noise. This requires either using a stronger laser or increasing the exposure time. A stronger laser can cause heating, which may knock the atom out of its trap. Conversely, a longer exposure slows the readout and limits the rate at which quantum circuits can be executed.
Additionally, the fidelity of readout depends on many experimental parameters, including the numerical aperture of the imaging system, the qubit species, and spacing between atoms.
Effect of choice of atomic species on fast readout. Qubit readout remains a key bottleneck in neutral atom quantum computing, with significant variations in fidelity and speed across available atomic species, including Cs, Rb, Yb, and Sr illustrated in Figure 2. The performance of fluorescence-based state detection depends strongly on species-specific properties such as cycling transition strength, background scattering, and susceptibility to state leakage. Each atomic species imposes distinct system-level constraints such as photon budget, trap spacing, and detection optics, necessitating a careful balance between fidelity, speed, complexity, and scalability. 444We note that a readout time of is used in this work as an example of short duration measurements. In practice the lower limit on atom measurement time depends on many experimental parameters including the specific atom and atomic transitions used, optical detection system efficiency, camera sensitivity, and noise. Advanced techniques suggest that it should be possible to acquire high SNR images in (scott2025lasercoolingqubitmeasurements). Adding high performance image processing as described here has the potential to further reduce the effective measurement time.
@@ 2.3. Why does fast readout matter?
Readout is essential for extracting useful information from a quantum computer, its speed and reliability pose a major challenge for scaling up neutral atom systems and supporting fast feedback in quantum error correction. The execution time of a fault-tolerant quantum program can be coarsely approximated based on logical operations, QEC cycle repetitions and time taken by each QEC cycle, as,
Each component in this model encapsulates a key bottleneck to the scalable, fault-tolerant quantum computation,
Non-Clifford Gates (e.g., gates) are necessary for universal quantum computation. They are the slowest operations due to magic state distillation (gidney2025factor).
-Rounds represent the number of QEC iterations needed to suppress logical errors for a code of distance . As expected, larger values are required for higher physical error rates, increasing overall runtime.
QEC cycle time is the duration of one round of quantum error correction, where parity checks (syndromes) are generated to track errors and keep logical qubits free of errors.
@@
Atom loading and rearrangement overhead. In neutral atom platforms, building a defect-free 2D qubit lattice requires stochastic loading555This initialization process is slow, imaging intensive, and requires repeated fluorescence-based measurements to identify occupied sites and reposition atoms using optical tweezers. followed by iterative rearrangement, or near-deterministic loading from a reservoir which nevertheless requires a verification imaging step to achieve perfect array filling. (Norcia2024; Gyger2024). Fast, high-fidelity readout accelerates each stage of this pipeline—enabling denser lattices, reducing the probability of atom loss during relocation, and improving experiment yield. Though often hidden behind “setup time,” inefficiencies here can dominate runtime in large-scale systems or in hybrid pipelines where reloading is triggered adaptively.
QEC bootstrapping and syndrome verification. For QEC codes to start correctly, not only initialization, but ancilla qubits must be verified through measurement. Faster readout enables quicker feedback, reducing dead time and improving the throughput of syndrome generation across the code lattice.
Pipelining and reuse of physical infrastructure. In principle, multiple Magneto-Optical Traps (MOTs) with reloading tweezers (Norcia2024; Gyger2024) can be deployed in parallel to mask atom loading time, but doing so increases system complexity, cost, and physical footprint. Faster readout alleviates the need for deep pipelining by making each atom loading pass more productive, reducing the amortized cost per qubit and easing scalability constraints.
3. Understanding challenges in fast readout
Readout speed directly governs QEC cycle time, its influence is far broader—impacting lattice preparation, code startup, and system utilization efficiency. Fast readout is not just a lever for shorter execution times; it is a key enabler for practical, reliable, and reconfigurable fault-tolerant quantum computing on neutral atom platforms. In this work, we explore the advantages of employing machine learning to improve the accuracy and speed of measuring atomic qubits.